OpenStudy (anonymous):

can the quadratic formula solve 2x^2 - 5 = 0?

OpenStudy (anonymous):

Sure. 2x^2 = 5 x^2 = 5/2 x = +/- SQRT(5/2)

OpenStudy (anonymous):

The quadratic formula is completely general, it can solve any quadratic equation.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

So it's ok if b=0?